OpenStudy (anonymous):

The container that holds the water for the football team is 2/3 full. After pouring out 2 gallons of water, it is 1/2 full. How many gallons can the container hold?

OpenStudy (anonymous):

12 gallons

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

in case you were wondering, \[\frac{2}{3}-\frac{1}{2}=\frac{1}{6}\] so one sixth of the container is 2 gallons, so it holds 12